An office supply store recently sold a black printer ink cartridge for ​$18.99 and a color printer ink cartridge for ​$21.99. At the start of a recent fall​ semester, a total of 44 of these cartridges was sold for a total of ​$868.56. How many of each type were​ purchased?
Answer by Boreal(15235)

A+B=44, A=black and B=color
18.99A+21.99B=868.56
-18.99A-18.99B=-835.56
3B=33
B=11 color cartridges for $241.89
A=33 black cartridges for $626.67
They add to $868.56